2. Find and classify all critical points for f(x,y) = -22 + y? (x - 8)
Answer #1

frary) f(ay) -x²+y^(7-8) -x²+ ay²-84² critical points conditions af of oy 0 - 0 OY - 32+ y*=0 22= 9 ye tray -16y=0 2417 - 8)-(814) ot-s² -4(8) +32-4704) -36+ 36-64 --64 < 0 f has saddle point at (814) (81-4) = ot-5=-4 (8) + 32 -4 (-412 -32 738 -60